OpenStudy (anonymous):

If tan x degrees =z/10 and cos x degrees=10/y what is the value of sin x degrees? sin x = z/y sin x= y/z sin x=10z sin x=10y

myininaya (myininaya):

@grex10 try to use that tan(x)=sin(x)/cos(x)

OpenStudy (anonymous):

i got c using tan x = sin x/ cos x

myininaya (myininaya):

\[\sin(x)=\cos(x) \tan(x)\]

myininaya (myininaya):

now replace cos(x) with 10/y and tan(x) with z/10

myininaya (myininaya):

I just took tan(x)=sin(x)/cos(x) and multiply both sides by cos(x) to get that identity above I'm using

OpenStudy (anonymous):

so the answer is sin x = y/z?

myininaya (myininaya):

I think you are working through it too fast

OpenStudy (anonymous):

well its a timed test

myininaya (myininaya):

Oh I can't help you on a test.

OpenStudy (anonymous):

i only get 90 minutes

OpenStudy (anonymous):

yes u can

OpenStudy (anonymous):

ur no help at all. any one wanna help ?